OpenStudy (anonymous):

A(0,8) B(6,0) C(15,0) what's the area of the Triangle in square units?

OpenStudy (anonymous):

I think it is 36 sq units

OpenStudy (anonymous):

how did you get that answer tho... im confused

OpenStudy (anonymous):

well we use the formula (1/2)* base * height we take BC as the base.... since both points are on the x axis the distance between the points (base length) is 15 - 6 = 9 units # the height is the perpendicular distance from base that is.. the distance of A from the x axis (y coordinate) which is 8 units so we get AREA OF TRIANGLE IS (1/2)*9*8 = 4*9 = 36 square units! get it now??

OpenStudy (anonymous):

ohhhh! now I get it.. sorta Thanks!

OpenStudy (anonymous):

so its the same thing as the perimeter... right?

OpenStudy (anonymous):

gimme a sec..

OpenStudy (anonymous):

YAAA YOU ARE RIGHT :P it is 36 again!! for finding the perimeter you've gotta use DISTANCE FORMULA finde distance between (A and B) then (B and C) then (C and A) then add em all up

OpenStudy (anonymous):

10 units + 9 units + 17 units = 36 units Good luck!!

OpenStudy (anonymous):

thank you!!

OpenStudy (anonymous):

You're most welcome ;)